OpenStudy (diamondanon2):

Perform the required operations on the following functions. Given: f(x) = 3 - x; g(x) = -2x Find f[g(2)].

OpenStudy (drmoss):

rly you gotta bump ahead lol

OpenStudy (drmoss):

im just kidding

OpenStudy (diamondanon2):

lol @jhonyy9

OpenStudy (drmoss):

?

OpenStudy (diamondanon2):

Answer choices are -1 7 25

OpenStudy (drmoss):

halp me ;(

jhonyy9 (jhonyy9):

so in this case you need finding f(g(2) first i say you can finding the g(2) = ? do you can ?

OpenStudy (diamondanon2):

-2(2)

jhonyy9 (jhonyy9):

yes and equal how many ?

OpenStudy (diamondanon2):

-4

jhonyy9 (jhonyy9):

exactly and now you just need to find f(-4) = ?

OpenStudy (diamondanon2):

3 * -4 = -12. But that is not an answer choice.

OpenStudy (diamondanon2):

or do we add

OpenStudy (diamondanon2):

as -4 + 3 = -1

jhonyy9 (jhonyy9):

f(x) = 3-x

OpenStudy (diamondanon2):

oh okay so 3 - (-4) = 7

jhonyy9 (jhonyy9):

exactly - very nice - this is the right answer sure hope you understand it now yes ?

OpenStudy (diamondanon2):

Yes thank you very much !

OpenStudy (drmoss):

can you help me after jhonny

jhonyy9 (jhonyy9):

do you have again ?

jhonyy9 (jhonyy9):

moment Moss - i will try after this - ok. ?

OpenStudy (drmoss):

ok :>

OpenStudy (drmoss):

she already closed this

jhonyy9 (jhonyy9):

@DiamondAnon2 do you have again ?